[Brahms-l] Shift report 20040316 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 16 2004 - 07:58:33 EST
------------------------------------------------------------
Date: Tue Mar 16 07:58:32 2004, Shift: 00:00-08:00
Supervisor: PS, with: 
Log book#15 pages 89 - 89

Shift Summary:
------------------------------------------------------------
1. beam through the most of the shift. Had two nice stores.

2. Collected 271k of tr 2 and 351k of tr 3 in runs
   from 11065 to 11083.

3. No detector problems detected.
   Problem with Io- and Run- Server still persists. 
    
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 16 07:59:47 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 16 2004 - 08:00:06 EST